simplify.(show your work)
- 7 - 8 + 5 + (-2) - 6

Answer by fractalier(2101)

Since we add and subtract as we find them left to right, we can do this by keeping a running total as we move left to right...thus
- 7 - 8 + 5 + (-2) - 6 =
-15 + 5 + (-2) - 6 =
-10 + (-2) - 6 =
-12 - 6 =
-18